OpenStudy (jdawg97):

who else here does Plato?? just curious (:

rebeccaxhawaii (rebeccaxhawaii):

post a specific question please. we will be happy to help.

OpenStudy (soprano.h.d0816):

I use Plato, what subject are you using it for?

ILovePuppiesLol (ilovepuppieslol):

Welcome to open study :)

OpenStudy (jdawg97):

i have to do English 12A and American Literature A and I'm struggling with both.. what are you doing? and thanks!

OpenStudy (soprano.h.d0816):

Errr... I do Biology and Geometry on there

OpenStudy (elonasushchik):

SPPPPPPPPPPPPPPASSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSMS

RhondaSommer (rhondasommer):

\(\Huge\color{limegreen}{\mathbb{WELCOME}}\) \(\Huge\color{forestgreen}{\mathbb{TO}}\) \(\Huge\color{limegreen}{\mathbb{OPENSTUDY}}\)

OpenStudy (just_one_last_goodbye):

Welcome to OpenStudy. I hope you enjoy your stay here. If you have any questions about the site ask the ambassadors. They have a purple Ⓐ by there name :) Or you could always shoot me a PM.

likeabossssssss (likeabossssssss):

^^